Subject: APO 3.5m Users committee minutes, 1/11/2010

From: Suzanne Hawley

Submitted: Wed, 20 Jan 2010 19:41:43 -0800 (PST)

Message number: 1167 (previous: 1166, next: 1168 up: Index)

       APO 3.5-m Users Committee Phonecon, 1/11/10

Attending:  Suzanne Hawley, Mark Klaene, Bill Ketzeback, Russet McMillan,
Al Harper, Bill Ketzeback, John Bally, Michael Strauss, Jon Holtzman, Scott 

Absent:  Jon Fulbright, Remy Indebetouw


User feedback, comments from institutional representatives:

Chicago (Al Harper) - nothing to report.

New Mexico State (Jon Holtzman) - nothing to report.

Washington (Scott Anderson) - thanks were offered to observers
and staff for helping with recent rescheduling of programs to
accomodate simultaneous Chandra observations.

Colorado (John Bally) - there was a good Triplespec run for CU recently, 
otherwise nothing to report.

Princeton - new postdoc Cullen Blake's recent trip to APO for training was very 
successful.  Also grateful for some recent
OPEN time assignments.

Johns Hopkins - no report (Fulbright absent)

Virginia - no report (Indebetouw absent)


Discussion of telescope/instruments report - Klaene:

Mark called attention to some of the failures and problems associated
with the recent cold weather, including the DIS slitviewer shutter
sticking (see report below).  There was also discussion of on-chip guiding with 
Agile.  Users committee members were encouraged to communicate with Agile users 
at their institutions, to let them know that on-chip guiding is being 
implemented soon in a
"quick and dirty" way that will likely require some special use
training.  A more extensive rewrite of the 3.5m guiding software is
planned during the next year that should make the Agile guiding
more robust.

Russet discussed the high altitude pointing and tracking issues.
Users should be aware that there are some disadvantages to following
objects through high altitude transits, including loss of tracking
and poor guiding.  In addition, it takes some time for the telescope
tracking to recover from such a transit, which involves taking the telescope to 
low elevation for reset.  The Observatory is still
recommending against high altitude observing (above 85 degrees),
and will only allow it in special circumstances.

Russet talked about the new Triplespec instrument block. There is a
great improvement to boresight guiding however field star guiding
will be much less accurate the farther from the slit you observe.
Users should talk to the Observing Specialist for advice if they
want to do field star guiding.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 12/4/09 through 

                          Mark Klaene

0) Overview

Weather, specifically snow, has been the main issue this period and we expect 
more in the coming months.  Large snow falls have created a lot of staff access 
problems as well as power losses.  During the first event we lost all site 
power for about 10 minutes as we scrambled to fix a generator problem. The good 
news is the generator is fixed and we now are carrying twice as much fuel on 
site. Loss time due to snow on the roof was minimal, however humidity did keep 
us closed on some occasions under clear skies. The holiday season came and went 
with many of the day staff on holiday leave but no major problems and only a 
couple minor issues during this time.

1) Telescope

Behavior has been quite good and a recent pointing model has us back to below 
5" RMS pointing.   Tracking improvements are still being looked into especially 
at high altitude. Recent TSpec instrument block update has also improved TSpec 
bore-sight guiding and nodding.

2) Instruments

DIS, Echelle, SPIcam, Agile and TripleSpec ran with no problems. DIS blue ION 
pump is running a little high.

The DIS slitviewer camera shutter is having its usual cold weather problems. We 
are monitoring this for increasing frequency.

The new NICFPS fanout board is progressing slowly with the board at CU for 
final assembly and testing.   Installation could be late January or February. 
There was no progress on the NIC-FPS etalon at CU.

3) CIF projects

The CIF projects for 2010 have been identified.  Major initiatives
include: calibration lamps (improved UV/blue flux intensity), high-altitude 
tracking, purchase and fabrication of direct-drive spares, and installation of 
a brake on the altitude axis.  Some money has been
set aside for instrument upgrades once a plan is in place for how to proceed.


o NICFPS fanout board, possible scheduling impact - Klaene

John Bally had no update on the status of the fanout board testing
at Colorado. In December, we were hopeful that we would get a fully populated 
and tested fanout board delivered to APO by the end of January, which would 
then be installed during a block of time when the instrument is not in use.  It 
is not clear if we
will be able to meet this schedule.  Thermal cycling of the
instrument, opening it up and installing the board will take
around two weeks.  Users should be aware that there may be some
impact on science operations if there are problems during the
installation process. If the fanout board replacement is unsuccessful
on the mountain, the instrument may have to go back to Colorado for
completion of work.

o Triplespec scheduling in Q1 - Hawley

Triplespec is a heavy, awkward instrument that is particularly difficult to 
mount and dismount in cold weather.  Therefore,
we are avoiding instrument changes that involve mounting or
dismounting Triplespec at midnight during Q1.  Users should
keep this restriction in mind when requesting OPEN or DD time
during Q1.  We will reevaluate this operational restriction
for Q2.

o FIRST/echelle/imager updates - Hawley

FIRST/IRET: There was a meeting with Jian Ge about the FIRST/IRET instrument in 
early December.  We asked for a fixed optical design prior to PDR, which Ge has 
promised to provide.  The project is on hold until the optical design is 

Echelle upgrade:  We are presently searching for a PI for
the echelle upgrade project, with the intention of proceeding
with this project in the near future.  There have been some concerns
expressed about the echelle availability.  With our current budget and schedule 
expectations, significant downtime is not likely for
at least 18 months.

Imager upgrade:  This project is on hold for now, pending
additional investigation into design options.



[open from previous months]: NONE

[new actions from this meeting]: NONE


Next meeting: The next Users Committee phonecon will be on Monday, 8 February, 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent to the 
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1167 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to